# Meta-Learning Objectives for Preference Optimization

### v2

@@ 5 Experiments: LLM transfer
We show that the insights obtained in the MuJoCo environments can be transferred to the LLM alignment setting. In particular, we found that all our discovered loss functions kept optimizing the policy even when $\pi(\tau_{w})>\pi(\tau_{l})$, while all algorithms in Table 1 have a PO term whose gradient quickly reaches 0 when $\pi(\tau_{w})>\pi(\tau_{l})$.
Using this insight, we design a new PO objective, which we call Constrained Linear Preference Optimization (CoLPO) and define as
$\pi^{\star}\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}\big{[}\log\pi_{\theta}(\tau_{w}|x)$
$+\beta\mathbf{1}_{\{\log\pi_{\theta}(\tau_{w}|x)<\alpha\}}\min(\log(\pi_{\theta}(\tau_{w}))-\log(\pi_{\theta}(\tau_{l})),0)$
$+\beta\mathbf{1}_{\{\log\pi_{\theta}(\tau_{w}|x)>\alpha\}}\min(\log(\pi_{\theta}(\tau_{w}))-\log(\pi_{\theta}(\tau_{l})),\gamma)\big{]},$
where $\beta\geq 0$, $\gamma\geq 0$ and $\alpha\leq 0$ are hyperparameters. The CoLPO objective keeps optimizing the policy when $\pi(\tau_{w})>\pi(\tau_{l})$, but only if $\log\pi_{\theta}(\tau_{w}|x)$ is over a certain threshold. This constraint is placed to stop the optimizer from increasing the difference $\log(\pi_{\theta}(\tau_{w}))-\log(\pi_{\theta}(\tau_{l}))$ by descreasing both probabilities.
We evaluate CoLPO on an LLM fine-tuning task, where we fine-tune the gemma-7b (Team et al., 2024) model on the dpo-mix-7k [^2] https://huggingface.co/datasets/argilla/dpo-mix-7k dataset for 3 epochs. Being a real-world preference dataset, dpo-mix-7k naturally presents characteristics of the mixed-quality and the noisy settings. To tune the LLMs, we modify the Alignment Handbook library (Tunstall et al., ) to include the CoLPO objective. We evaluate the tuned LLM against GPT-4, using the AlpacaEval library  (Li et al., 2023). All models were trained using the default hyper-parameters, with $\beta=0.5$, $\gamma=5$ (for CoLPO and SLiC-HF), and $\alpha=-1$ (for CoLPO). For details on hyperparameters and training settings consult Appendix J.
Table 3 reports the results of CoLPO, as well as the results for other one-step algorithms, and shows that CoLPO largely outperforms existing algorithms, confirming the validity of the insights gained in the MuJoCo simulations.
Lastly, we report that the the log-probabilities for the final training batch of chosen and rejected responses are -1.17 and -11.38 respectively for CoLPO, compared to -0.61 and -2.98 for ORPO, respectively.
This results further confirms that CoLPO is effective in decreasing the probability of rejected responses and mantaining high the probability of chosen responses.
@@ 6 Conclusion
We have introduced a novel framework for Preference Optimization algorithms, as well as a methodology for the automatic discovery of PO algorithms using evolutionary strategies. Through a systematic evaluation across diverse settings in MuJoCo environments, we have demonstrated that the performance of our discovered objectives consistently exceeds the performance of existing methods, particularly in noisy and mixed-quality datasets where many baselines underperform.
Our analysis in MuJoCo also revealed a common shortcoming among current baselines: truncating the loss whenever $\pi(\tau_{w})>\pi(\tau_{l})$. To address this, we proposed a simple objective, CoLPO, that avoids such loss truncation while preventing the optimizer from decreasing the probability of chosen datapoints. We then tested this objective on an LLM fine-tuning task, achieving significant improvements over existing methods, thereby confirming the broader applicability of our approach.
Appendix A Popular PO Algorithms
These methods include RRHF, which uses length-normalized log-likelihood, and SLiC-HF, which uses direct log-likelihood and incorporates SFT.
The comparison also includes IPO, a theoretically-based approach that handles pairwise preferences differently than DPO. Another method, CPO, combines sequence likelihood as a reward with an SFT objective. Finally, R-DPO modifies the original DPO by adding regularization to prevent length exploitation.
@@ Appendix B Related Work
Several works in the literature have shown that it is possible to discover machine learning algorithms that outperform algorithms manually designed by researchers (Oh et al., 2020; Lu et al., 2022; Jackson et al., 2024; Alfano et al., 2024). An approach particularly relevant to our method is DiscoPOP by (Lu et al., 2024), which leverages an LLM to discover objective functions for LLM tuning. They consider a different space of objective functions from us, as they replace the log-sigmoid in (5) with a generic loss function, following the framework built by (Tang et al., 2024). Additionally, instead of searching over a space of parametrized functions, they ask the LLM to generate loss functions in code space. This distinction suggests that our approaches could be complementary, as the model discovered by DiscoPOP could be paired with our learned mirror map. Lastly, DiscoPOP optimizes its objective function directly on the final task, whereas we adopt a two-stage process—optimizing the loss function on a separate task (MuJoCo) and later transferring it to the LLM setting. This transferability underscores the broader applicability of our approach.
@@
A generalization of DPO alternative to ours is $f$-DPO, developed by (Wang et al., 2023), which consists in replacing the $\mathtt{KL}$-divergence in (1) with an $f$-divergence and then apply the same heuristic as DPO to obtain the final objective function. We note that the $\mathtt{KL}$-divergence is the only $f$-divergence to be also a Bregman divergence, and vice-versa. They empirically demonstrate that different $f$-divergences lead to different balances between alignment performance and generation diversity, highlighting the trade-offs inherent to this class of algorithms. (Huang et al., 2024) further explore this class of PO algorithm and individuate an $f$-divergence for which $f$-DPO is robust to overoptimization.
@@
We provide here a proof for our main result, i.e. Theorem 3.1. The proof to obtain the DPO objective in (5) follows by taking $\phi=e^{x}$.
Theorem C.1 (Theorem 3.1).
Let $h_{\phi}$ be a 00-potential mirror map and $\pi^{\star}$ be a solution to the optimization problem in (7). If $\pi_{\mathrm{ref}}(a|s)>0$ for all $s\in\mathcal{S},a\in\mathcal{A}$, we have that
@@
$\nabla_{\pi(\tau)}$ $\Bigg{[}\sum_{t=0}^{T-1}\mathbb{E}_{a\sim\pi(\cdot|s_{t})}[r(s_{t},a)]-\beta D_{h}(\pi(\cdot|\tau),\pi_{\mathrm{ref}}(\cdot|\tau))$
$\quad-\lambda\left(\sum_{\tau^{\prime}:s_{0}\in\tau^{\prime}}\pi(\tau^{\prime})-1\right)+\sum_{\tau^{\prime}:s_{0}\in\tau^{\prime}}\alpha(\tau^{\prime})\pi(\tau^{\prime})\Bigg{]}$
$=r(\tau)-\beta\phi^{-1}(\pi(\tau))+\phi^{-1}(\pi_{\mathrm{ref}}(\tau))-\lambda+\alpha(\tau)=0,$
@@
$\pi(\tau)=\phi\big{(}r(\tau)+\phi^{-1}(\pi_{\mathrm{ref}}(\tau))-\lambda+\alpha(\tau)\big{)}.$
Since $0\notin\mathop{\mathrm{dom}}\phi^{-1}$, due to the definition of a 00-potential, and $\pi_{\mathrm{ref}}(\tau)>0$, we have that $\pi(\tau)>0$ for all trajectories $\tau$. Invoking the complementary slackness condition, whereby $\alpha(\tau)\pi(\tau)=0$ for all trajectories $\tau$, we have that $\alpha(\tau)=0$ for all trajectories $\tau$. Therefore, we have that
$r(\tau)-\beta\phi^{-1}(\pi(\tau))+\phi^{-1}(\pi_{\mathrm{ref}}(\tau))-\lambda=0$
@@
Appendix D Further discussion of $\omega$-potentials
@@
$\mathcal{D}_{h}(x,y)$ $=h(x)-h(y)-\langle\nabla h(y),x-y\rangle$
$=\sum_{a\in\mathcal{A}}x_{a}^{2}-y_{a}^{2}-(2y_{a})(x_{a}-y_{a})$
$=\sum_{a\in\mathcal{A}}(x_{a}-y_{a})^{2},$
@@
Appendix E Further discussion on Evolution Strategies
@@
In our setting, we use ES to search for the best $\psi$ and $\phi^{-1}$ within the parametrized class introduced in Section 3.1, so that an agent trained using the objective in (10) achieves the highest value. Denote by $\zeta$ the parameters of $\psi$ and $\phi^{-1}$ and by $\pi^{\zeta}$ the final policy obtained optimizing the objective in (10) when using the parametrized $\psi$ and $\phi^{-1}$. Lastly, let $F(\zeta)$ be the expected cumulative reward (or value) of $\pi^{\zeta}$, i.e. $F(\zeta)=\mathbb{E}_{\tau\sim(\mu,\pi^{\zeta},P)}r(\tau)$. At each iteration, we estimate the gradient $\nabla_{\zeta}F(\zeta)$ as
@@
Appendix F Further discussion on Online vs Offline Methods

### v3

@@ 5 Experiments: LLM transfer
We show that the insights obtained in the MuJoCo environments can be transferred to the LLM alignment setting. In particular, we test the TA-MPO algorithm in (13), which is designed to continue to optimize the policy even when $\pi(\tau_{w})>\pi(\tau_{l})$, as all our discovered algorithms do. We evaluate TA-MPO on LLM alignment, comparing its performance against baselines for three combinations of base model and preference dataset, as shown in Table˜2.
To tune the LLMs, we modify the Alignment Handbook library (Tunstall et al., ) to include the TA-MPO objective in (13). We evaluate the tuned LLMs against GPT-4, using the AlpacaEval library  (Li et al., 2023) and Llama-3.1-70B-Instruct as a judge. For all combinations of starting LLM, dataset, and PO algorithm, we perform 4 update epochs and set the learning rate to 5e-5 and $\beta$ to 0.05. In the case of DPO and SimPO, we performed 3 epochs of SFT, with learning rate 5e-5, and 1 epoch of DPO/SimPO, with learning rate 5e-7 and $\beta=0.05$. Refer to Appendix˜L for further details.
Table˜2 shows the effectiveness of the TA-MPO objective in (13), which presents a high winrate for all schedules of $\alpha$. In particular, TA-MPO with the sigmoid schedule for $\alpha$ has the highest winrate in all tasks and the highest length controlled winrate in two out of three tasks. Additionally, TA-MPO requires only one stage of training, while DPO and SimPO require two, i.e. SFT and PO.
We also tested the static algorithms discovered in the MuJoCo environment. Those with $g=\log\sigma$, which rediscovered CPO, exhibited performance similar to CPO itself. In contrast, algorithms with a parameterized $g$ function, which learned to greedily optimize the policy even when $\pi(\tau_{w})>\pi(\tau_{l})$, performed poorly. We hypothesize that these algorithms overfit to the TLA task, where our datasets sufficiently cover the state-action space. On the other hand, in LLM tuning—where data is sparse relative to the environment—greedy methods are more susceptible to over-optimization. At the same time, the objectives discovered on TLA are only used to logits within -14 and -8, as shown in Figure˜6, and have less regular shape outside of this subset.
In Table˜3, we report additional results that allow us to better understand the influence of different components of TA-MPO. In particular, we have tested SimPO with $\gamma=0$, SimPO with $\gamma>0$, and SFT + SimPO with $\gamma>0$, where the last one consists of a one-step objective made of the addition between the SFT and the SimPO loss. Table˜3 shows that having a large $\gamma$, which encourages large updates even when $\pi(\tau_{w})>\pi(\tau_{l})$, or adding an SFT term to the PO loss are both helpful in improving performance. The performance of TA-MPO reported in Table˜2 is still the highest, which means that temporal awareness also contributes to a better performance. We can therefore confirm that each of the insights we gained in the MuJoCo benchmark transfers to the LLM alignment setting.
Table 3: We report win-rates and standard error on AlpacaEval (length controlled win-rates and standard error in parenthesis) for gemma-7b and two preference dataset.
Algorithm $\gamma$ dpo-mix-7k capybara-7k
SimPO 0 30.50 (26.06) 27.52 (25.12)
SimPO 1 30.62 (25.35) 27.52 (25.54)
SimPO 2 32.50 (25.80) 27.95 (25.71)
SimPO 5 33.42 (28.22) 27.02 (25.07)
SimPO 10 33.60 (27.12) 27.39 (25.00)
SFT+SimPO 1 30.31 (26.55) 35.07 (28.25)
@@ 6 Conclusion
We have introduced a novel framework for Preference Optimization algorithms, as well as a methodology for the automatic discovery of PO algorithms using evolutionary strategies. Through a systematic evaluation across diverse settings in MuJoCo environments, we have demonstrated that the performance of our discovered objectives consistently exceeds the performance of existing methods, particularly in noisy and mixed-quality datasets where many baselines underperform. Our analysis in MuJoCo also revealed a common shortcoming among current baselines: truncating the loss whenever $\pi(\tau_{w})>\pi(\tau_{l})$. Using this insight, we proposed a temporally-aware algorithm, TA-MPO, that avoids such loss truncation and gradually switches from the SFT step to the PO step. We then tested this objective on an LLM fine-tuning task, achieving significant improvements over existing methods, thereby confirming the broader applicability of our approach.
Acknowledgments and Disclosure of Funding
Carlo Alfano and Silvia Sapora are supported by the Engineering and Physical Sciences Research Council EP/W524311/. Jakob Foerster is partially funded by the UKI grant EP/Y028481/1 (originally selected for funding by the ERC). Jakob Foerster is also supported by the JPMC Research Award and the Amazon Research Award. Patrick Rebeschini is funded by UK Research and Innovation (UKRI) under the UK government’s Horizon Europe funding guarantee [grant number EP/Y028333/1]. Yee Whye Teh acknowledges support from the Ministry of Digital Development and Information (MDDI) under the Singapore Global AI Visiting Professorship Program (Award No. AIVP-2024-002).
Appendix A Baseline PO Algorithms
Table 4: Overview of popular PO algorithms. The objective is to be maximized and $(\tau_{w},\tau_{l})\sim\mathcal{D}$.
Method Objective
IPO (Azar et al., 2024) $\left(\log\frac{\pi_{\theta}(\tau_{w}|x)}{\pi_{\mathrm{ref}}(\tau_{w}|x)}-\log\frac{\pi_{\theta}(\tau_{l}|x)}{\pi_{\mathrm{ref}}(\tau_{l}|x)}-\frac{1}{2\tau}\right)^{2}$
These methods include RRHF, which uses length-normalized log-likelihood, and SLiC-HF, which uses direct log-likelihood and incorporates SFT. The comparison also includes IPO, a theoretically-based approach that handles pairwise preferences differently than DPO. Another method, CPO, combines sequence likelihood as a reward with an SFT objective. Finally, R-DPO modifies the original DPO by adding regularization to prevent length exploitation.
@@ Appendix B Related Work
Several works in the literature have shown that it is possible to discover machine learning algorithms that outperform algorithms manually designed by researchers (Oh et al., 2020; Lu et al., 2022; Jackson et al., 2024; Alfano et al., 2024). An approach particularly relevant to our method is DiscoPOP by Lu et al. (2024), which leverages an LLM to discover objective functions for LLM tuning. They consider a different space of objective functions from us, as they replace the log-sigmoid in (5) with a generic loss function, following the framework built by Tang et al. (2024). Additionally, instead of searching over a space of parametrized functions, they ask the LLM to generate loss functions in code space. This distinction suggests that our approaches could be complementary, as the model discovered by DiscoPOP could be paired with our learned mirror map. Lastly, DiscoPOP optimizes its objective function directly on the final task, whereas we adopt a two-stage process—optimizing the loss function on a separate task (MuJoCo) and later transferring it to the LLM setting. This transferability underscores the broader applicability of our approach.
@@
A generalization of DPO alternative to ours is $f$-DPO (Wang et al., 2023), which consists in replacing the $\mathtt{KL}$-divergence in (1) with an $f$-divergence and then apply the same heuristic as DPO to obtain the final objective function. We note that the $\mathtt{KL}$-divergence is the only $f$-divergence to be also a Bregman divergence, and vice-versa. They empirically demonstrate that different $f$-divergences lead to different balances between alignment performance and generation diversity, highlighting the trade-offs inherent to this class of algorithms. Huang et al. (2024) further explore this class of PO algorithm and individuate an $f$-divergence for which $f$-DPO is robust to overoptimization.
@@
We provide here a proof for our main result, i.e. Theorem˜3.1. The proof to obtain the DPO objective in (5) follows by taking $\phi=e^{x}$.
Theorem C.1 (Theorem˜3.1).
Let $h_{\phi}$ be a 0-potential mirror map and $\pi^{\star}$ be a solution to the optimization problem in (7). If $\pi_{\mathrm{ref}}(a|s)>0$ for all $s\in\mathcal{S},a\in\mathcal{A}$, we have that
@@
$\nabla_{\pi(\tau)}$ $\Bigg[\sum_{t=0}^{T-1}\mathbb{E}_{a\sim\pi(\cdot|s_{t})}[r(s_{t},a)]-\beta D_{h}(\pi(\cdot|\tau),\pi_{\mathrm{ref}}(\cdot|\tau))-\lambda\hskip-8.5359pt\sum_{\tau^{\prime}:s_{0}\in\tau^{\prime}}\hskip-8.5359pt\pi(\tau^{\prime})-\lambda+\hskip-8.5359pt\sum_{\tau^{\prime}:s_{0}\in\tau^{\prime}}\hskip-8.5359pt\alpha(\tau^{\prime})\pi(\tau^{\prime})\Bigg]$
$=r(\tau)-\beta\phi^{-1}(\pi(\tau))+\beta\phi^{-1}(\pi_{\mathrm{ref}}(\tau))-\lambda+\alpha(\tau)=0,$
@@
$\pi(\tau)=\phi\big((r(\tau)+\beta\phi^{-1}(\pi_{\mathrm{ref}}(\tau))-\lambda+\alpha(\tau))/\beta\big).$
Since $0\notin\mathop{\mathrm{dom}}\phi^{-1}$, due to the definition of a 0-potential, and $\pi_{\mathrm{ref}}(\tau)>0$, we have that $\pi(\tau)>0$ for all trajectories $\tau$. Invoking the complementary slackness condition, whereby $\alpha(\tau)\pi(\tau)=0$ for all trajectories $\tau$, we have that $\alpha(\tau)=0$ for all trajectories $\tau$. Therefore, we have that
$r(\tau)-\beta\phi^{-1}(\pi(\tau))+\beta\phi^{-1}(\pi_{\mathrm{ref}}(\tau))-\lambda=0$
@@
Appendix D Reward Modeling
In Equation (9), we utilize the interpretation of reward modeling as a binary classification problem given by Tang et al. (2024), which we summarize here. Let $z=(\tau_{1},\tau_{2})$ be a pair of trajectories and $\ell\in\{-1,1\}$ be the associated label that states whether $\tau_{1}$ is preferred to $\tau_{2}$ ($\ell=1$) or not ($\ell=-1$). We want to find a function $\hat{\ell}(z)\in\mathbb{R}$ such that $\mathop{\mathrm{sign}}(\hat{\ell}(z))$ is a good estimate of $\ell$. For a dataset $\mathcal{D}=\{z_{i},\ell_{i}\}_{i=1}^{N}$, the classification loss (or 0-1 loss) is
$L(\widehat{\ell},\mathcal{D})=\mathbb{E}_{\mathcal{D}}\left[1-\text{sign}\left(\hat{\ell}(z)\cdot\ell\right)\right],$ (15)
which is often approximated with a surrogate
$L_{f}(\widehat{\ell},\mathcal{D})=\mathbb{E}_{\mathcal{D}}\left[f\left(\hat{\ell}(z)\cdot\ell\right)\right],$ (16)
for a function $f:\mathbb{R}\to\mathbb{R}$. This approximation is possible because, when $f$ is decreasing (or convex), (15) and (16) have the same minimizer, as we prove in the following. Denote $p_{1}(z)=\mathbb{P}(\ell=1|z)$, then the conditional surrogate loss at $z$ is
$L_{f}(\widehat{\ell},x)=p_{1}(z)f\left(\hat{\ell}(z)\right)+(1-p_{1}(z))f\left(-\hat{\ell}(z)\right).$ (17)
The minimizer $\widehat{\ell}^{\star}$ of (16) is such that $\widehat{\ell}^{\star}(z)$ minimizes (17). While the minimizer to (17) might not be computable explicitly, we can show $\widehat{\ell}^{\star}(z)>0\iff p_{1}(z)>1/2$ when $f$ is a decreasing function, meaning that $\widehat{\ell}^{\star}$ is also the minimizer of (15). Firstly, we have that
$L_{f}(\widehat{\ell},x)-L_{f}(-\widehat{\ell},x)$ $=p_{1}(z)f\left(\hat{\ell}(z)\right)+(1-p_{1}(z))f\left(-\hat{\ell}(z)\right)$ (18)
$\quad-p_{1}(z)f\left(-\hat{\ell}(z)\right)-(1-p_{1}(z))f\left(\hat{\ell}(z)\right)$
$=\left(2p_{1}(z)-1\right)\left(f\left(\hat{\ell}(z)\right)-f\left(-\hat{\ell}(z)\right)\right)$
Since $f$ is decreasing, we have for all $\widehat{\ell}(z)>0$ that $f(\widehat{\ell}(z))<f(-\widehat{\ell}(z))$. Plugging this into (18), we obtain that, if $p_{1}(z)>1/2$, $L_{f}(\widehat{\ell},x)<L_{f}(-\widehat{\ell},x)$ for all $\widehat{\ell}(z)>0$. Therefore, the minimizer $\widehat{\ell}^{\star}(z)$ of (17) must be positive. The opposite can be proved in the same manner.
Equation (9) can be obtained by setting f=gf=-g, for an increasing function $g$, and
$\widehat{\ell}(z)=\widehat{r}(\tau_{1})-\widehat{r}(\tau_{2}).$
Appendix E Bregman divergences vs $f$-divergences
In this section, we discuss why we use Bregman divergences rather than $f$-divergences in (7). Firstly, we note that, if the reference policy $\pi_{\mathrm{ref}}$ is set to be the uniform distribution, Bregman divergences and $f$-divergences generate equivalent families of algorithms. Let $\phi$ be a 0-potential. Using a reasoning similar to that of Theorem˜3.1, we have the following cases:
Bregman divergence $\quad r(\tau_{w})-r(\tau_{l})=\phi^{-1}(\pi^{*}(\tau_{w}))-\phi^{-1}(\pi_{\mathrm{ref}}(\tau_{w}))-\phi^{-1}(\pi^{*}(\tau_{l}))+\phi^{-1}(\pi_{\mathrm{ref}}(\tau_{l})),$
${\scriptsize\text{+uniform }\pi_{\mathrm{ref}}}$ $\quad r(\tau_{w})-r(\tau_{l})=\phi^{-1}(\pi^{*}(\tau_{w}))-\phi^{-1}(\pi^{*}(\tau_{l})),$
$f$-divergence $\quad r(\tau_{w})-r(\tau_{l})=\phi^{-1}(\pi^{*}(\tau_{w})/\pi_{\mathrm{ref}}(\tau_{w}))-\phi^{-1}(\pi^{*}(\tau_{l})/\pi_{\mathrm{ref}}(\tau_{l})),$
${\scriptsize\text{+uniform }\pi_{\mathrm{ref}}}$ $\quad r(\tau_{w})-r(\tau_{l})=\phi^{-1}(|\mathcal{T}|(\pi^{*}(\tau_{w})))-\phi^{-1}(|\mathcal{T}|(\pi^{*}(\tau_{l})))$
$\qquad\qquad\qquad\;\;=\tilde{\phi}^{-1}(\pi^{*}(\tau_{w}))-\tilde{\phi}^{-1}(\pi^{*}(\tau_{l})),$
where $\mathcal{T}$ is the set of all trajectories and $\tilde{\phi}^{-1}(x)=\phi^{-1}(|\mathcal{T}|x)$. The two resulting expressions are equivalent, meaning that our 1S-MPO class also includes the case of $f$-divergences. Our experiment on the 1S-MPO class therefore explore both Bregman divergences and $f$-divergences with ES and Table˜1 reports the performance of the best divergence found within both divergence classes.
If the reference policy is not uniform, then the two algorithmic classes have the KL divergence as the only intersection. In this setting, we have chosen to focus on Bregman divergences as Wang et al. (2023) have already shown that, among the $f$-divergences they considered, the KL-divergence typically offers superior alignment performance. Since we wanted to explore a class of algorithms with the objective of finding better alternatives to the KL-divergence, we decided to consider a different generalization of the KL-divergence, i.e. Bregman divergences. In our experiments with $g=\log$, we discovered that the KL-divergence is optimal also within Bregman divergences. On the other hand, we found out that significant improvements in performance come from modifying $g$ rather than the KL-divergence. We have observed similar results in preliminary MuJoCo experiments where we replaced the Bregman divergence with an $f$-divergence, that is that the KL-divergence is optimal among $f$-divergences.
Appendix F Further discussion of $\omega$-potentials
@@
$\mathcal{D}_{h}(x,y)$ $=h(x)-h(y)-\langle\nabla h(y),x-y\rangle=\sum_{a\in\mathcal{A}}x_{a}^{2}-y_{a}^{2}-(2y_{a})(x_{a}-y_{a})=\sum_{a\in\mathcal{A}}(x_{a}-y_{a})^{2},$
@@
Appendix G Further discussion on Evolution Strategies
@@
In our setting, we use ES to search for the best $\psi$ and $\phi^{-1}$ within the parametrized class introduced in Section˜3.1, so that an agent trained using the objective in (11) achieves the highest value. Denote by $\zeta$ the parameters of $\psi$ and $\phi^{-1}$ and by $\pi^{\zeta}$ the final policy obtained optimizing the objective in (11) when using the parametrized $\psi$ and $\phi^{-1}$. Lastly, let $F(\zeta)$ be the expected cumulative reward (or value) of $\pi^{\zeta}$, i.e. $F(\zeta)=\mathbb{E}_{\tau\sim(\mu,\pi^{\zeta},P)}r(\tau)$. At each iteration, we estimate the gradient $\nabla_{\zeta}F(\zeta)$ as
@@
Appendix H Further discussion on Online vs Offline Methods